### v1

@@ Abstract
3D Gaussian Splatting (3DGS) creates a radiance field consisting of 3D Gaussians to represent a scene. With sparse training views, 3DGS easily suffers from overfitting, negatively impacting the reconstruction quality. This paper introduces a new co-regularization perspective for improving sparse-view 3DGS. When training two 3D Gaussian radiance fields with the same sparse views of a scene, we observe that the two radiance fields exhibit point disagreement and rendering disagreement that can unsupervisedly predict reconstruction quality, stemming from the sampling implementation in densification. We further quantify the point disagreement and rendering disagreement by evaluating the registration between Gaussians’ point representations and calculating differences in their rendered pixels. The empirical study demonstrates the negative correlation between the two disagreements and accurate reconstruction, which allows us to identify inaccurate reconstruction without accessing ground-truth information. Based on the study, we propose CoR-GS, which identifies and suppresses inaccurate reconstruction based on the two disagreements: (i) Co-pruning considers Gaussians that exhibit high point disagreement in inaccurate positions and prunes them. (ii) Pseudo-view co-regularization considers pixels that exhibit high rendering disagreement are inaccurately rendered and suppress the disagreement. Results on LLFF, Mip-NeRF360, DTU, and Blender demonstrate that CoR-GS effectively regularizes the scene geometry, reconstructs the compact representations, and achieves state-of-the-art novel view synthesis quality under sparse training views. Project page: https://jiaw-z.github.io/CoR-GS/
@@
radiance fields 3d gaussian splatting few-shot novel view synthesis co-regularization
@@ 1 Introduction
We hope the observations and discussions in this paper can stimulate further thinking on the behavior disagreement of 3D Gaussian radiance fields.
@@ 2 Related Work
Radiance Fields. Radiance Fields are employed for reconstructing 3D scenes and synthesizing novel views. Neural Radiance Fields (NeRFs) [24] have experienced significant advancements recently, which learn the neural volumetric representations to represent 3D scenes and render images through volume rendering. Since then, a great number of studies aim to improve the rendering quality [2, 3] and efficiency [5, 25, 32, 11, 43, 21, 13] for NeRFs.
Recent advancements in unstructured radiance fields, as demonstrated by studies like [7, 41, 16], employ a collection of Gaussians to effectively represent 3D scenes. Notably, 3D Gaussian Splatting (3DGS) [16] adopts an approach where radiance fields are represented using a set of anisotropic 3D Gaussians, and renders images via differentiable splitting. 3DGS has demonstrated remarkable success in reconstructing high-quality complex scenes in real-time. Despite the tremendous success of 3DGS in many 3D tasks [22, 38, 33], the behavior of 3DGS with sparse views is less investigated, posing an open problem in the field.
Novel View Synthesis with Sparse Views. Novel view synthesis aims to generate unseen views of objects or scenes from a set of given images [48, 1]. While given sparse views, many methods tend to reconstruct degraded scenes and render unrealistic novel views. Currently, many studies have explored various regularization strategies for training NeRFs with sparse views [42, 26, 17, 10]. Some methods focus on designing generative models and pre-train them on extensive datasets [6, 44, 9, 49, 18]. Conversely, others [39, 14, 10, 27, 31, 35] the external knowledge from pre-trained models to regularize the training process.
With the advancements of 3DGS, some recent studies [50, 40, 19] have observed the issue of 3DGS to synthesize unrealistic novel views in sparse-view scenarios. These methods rely on predictions of pre-trained depth estimators as regularization to correct the reconstructed geometry. However, external supervision can introduce additional noise, negatively impacting the reconstruction. Our paper introduces a new co-regularization perspective for sparse-view 3DGS by suppressing the disagreement between two 3D Gaussian radiance fields.
Prediction Agreement. The agreement of predictions of two nerual networks has been explored in many tasks, especially for semi-supervised learning [30, 4, 20] and learning with noisy labels [12, 37, 45]. The typical pipeline involves simultaneously training two neural networks. They leverage the agreement of different networks’ predictions to pseudo-label unlabeled data [20] or clean noise from labeled annotations [12], thereby training networks with improved generalization capabilities. Our distinction from previous studies lies in our investigation of disagreement of unstructured 3D Gaussian radiance fields instead of generalized neural networks, where we consider Gaussian positions and rendered results.
@@ 3 Point Disagreement and Rendering Disagreement
In this section, we introduce the point disagreement and rendering disagreement, and demonstrate the negative correlation between the two disagreements and accurate reconstruction.
@@ 3.1 Preliminary and Definition
Point disagreement. When considering the 3-axis position of Gaussians, Gaussians in two radiance fields can be regarded as two point clouds. We evaluate their differences using Fitness and root mean square error (RMSE), which are used to evaluate the registration between point clouds [47]. Fitness calculates the overlapping area between Gaussians concerning a max points-pair distance $\tau=5$ for correspondence. RMSE computes the average distance of correspondent points.
@@ 3.2 Empirical Study
Two 3D Gaussian radiance fields trained with the same sparse views can exhibit different behaviors. We simultaneously train two 3D Gaussian radiance fields with the same sparse views and record the point disagreement and rendering disagreement during training. Since ground-truth supervision is directly imposed on training views, we evaluate the rendering disagreement at unseen views. As shown in Fig. 2, the two 3D Gaussian radiance fields exhibit different behaviors. In particular, the two disagreements grow significantly during densification, which creates new Gaussians but locates them blindly to the scene geometry. It also indicates that the optimization struggles to correct newly created Gaussians with sparse training views.
@@ 4.1 Co-pruning
Co-pruning leverages the point disagreement to identify Gaussians that are located in inaccurate positions. We perform co-pruning every certain number (we set it to 5) of the optimization/density control interleaves, as we aim to prune Gassuains that optimization cannot handle.
The positions of new Gaussians created by densification rely on subsequent optimization to correct their positions. However, under sparse constraints, if the positions of new Gaussians are in inaccurate locations, there is a probability that they cannot be corrected in subsequent optimization. This could lead to continued densification at inaccurate positions during scene reconstruction, ultimately resulting in incorrect geometric structures. Based on the point disagreement between two Gaussian radiance fields, co-pruning identifies and prunes Gaussians that are located in inaccurate positions. We first find the matching correspondence $f:\Theta^{1}\to\Theta^{2}$ based on their nearest Gaussians in the other set:
@@
where $\theta^{1}_{xi},\theta^{1}_{yi},\theta^{1}_{zi}$ represent the 3-axis positions of Gaussians. After the calculation of the non-matching mask for both sets of Gaussians, we prune all Gaussians that are marked as non-matching, which are considered in inaccurate positions.
@@ 4.2 Pseudo-view Co-regularization
Pseudo-view co-regularization leverages the rendering disagreement and conducts co-regularization on the pseudo-view rendering. The pseudo-view co-regularization is incorporated from the first densification step.
@@
For rendered depth maps, we find that although depth co-regularization alone helps, the impact becomes very weak when used in conjunction with color co-regularization, which is shown in the Supplementary Material. This is because the rendering process of colors also reflects the depth information of Gaussians due to the sorting process implied for rendering. Therefore, color co-regularization also imposes constraints on depth information and we do not perform co-regularization on depth maps in this case. The final training loss is a combination of the ground truth supervised loss at the training view and the color co-regularization term at the pseudo view:
@@ 5.1 Setup
Datasets. We conducted experiments on four datasets: the LLFF dataset [23], the Mip-NeRF360 dataset [3], the DTU dataset [15], and the NeRF Blender Synthetic dataset (Blender) [24]. Our experimental setup aligns with previous studies [26, 42, 35, 50], adopting the same data split, with downsampling rates of $8$, $4$, $4$, and $2$ for LLFF, Mip-NeRF360, DTU, and Blender, respectively. To focus on the target object and eliminate background noise during evaluation, we apply object masks similar to prior works [26] for DTU. Camera poses are assumed to be known based on calibration or other established methods, following the conventions of sparse-view settings.
@@
Implementations. Our implementation is based on the official PyTorch 3D Gaussian Splatting codebase. We conduct training for 10k iterations on the LLFF, DTU, and Blender datasets and follow 3DGS [16] to train models for 30k iterations on the MipNeRF dataset. In all experiments, we set the balance weight $\lambda_{p}=1.0$ for pseudo views in the losses. Following FSGS [50], we initialize 3DGS and CoR-GS with fused stereo point cloud from sparse views.
@@ 5.2 Comparison
LLFF. We provide the quantitative results on the LLFF dataset in Tab. 1. Our method consistently achieves the best performance across PSNR, SSIM, LPIPS, and AVGE metrics with 3, 6, and 9 training views. 3DGS-based methods exhibit a significant advantage in terms of scene structure compared to NeRF-based methods. The depth-supervised FSGS enhances the performance of 3DGS in scenarios with 3 and 6 views. However, with 9 training views, 3DGS can already reconstruct the scene structure well, which limits the effectiveness of external depth supervision. Our method conducts co-regularization to avoid reconstructing incorrect geometry, leading to improvements across all metrics with various training views. We provide quantitative visualizations of rendered images and depth at novel views and Gaussian points in Fig. 6. From the rendered depth maps, we observe that vanilla 3DGS can recover certain scene structures but exhibits geometric errors, particularly evident in the unrealistic parts of rendered images at novel views. For the point representations, many Gaussians are dispersed throughout space and far from the reconstructed scene. These Gaussians, when observed from novel views, can result in unrealistic rendering images. With depth supervision, FSGS can correct the incorrect geometry of 3DGS; however, noises in external depth maps also adversely affect the geometry. Similarly, relying on depth supervision makes it challenging to directly constrain Gaussians. It can be seen that the Gaussians’ dispersion of FSGS is still not compact. Our method effectively assists Gaussians in reconstructing coherent and compact geometry, resulting in higher-quality novel view rendering.
@@ 5.3 Ablation Study
Co-pruning. Co-pruning improves the reconstruction of the scene’s geometric structure, reflected in more reasonable depth rendering and more compact Gaussians. However, its drawback is that the rendered depth lacks smoothness, especially evident in the floor part of the horns scene. This is because co-pruning mainly operates on the non-matching Gaussians that are far from reconstructed scenes while lacking sufficient constraints on the matching Gaussians that are located in reasonable positions.

### v2

@@ Abstract
3D Gaussian Splatting (3DGS) creates a radiance field consisting of 3D Gaussians to represent a scene. With sparse training views, 3DGS easily suffers from overfitting, negatively impacting rendering. This paper introduces a new co-regularization perspective for improving sparse-view 3DGS. When training two 3D Gaussian radiance fields, we observe that the two radiance fields exhibit point disagreement and rendering disagreement that can unsupervisedly predict reconstruction quality, stemming from the randomness of densification implementation. We further quantify the two disagreements and demonstrate the negative correlation between them and accurate reconstruction, which allows us to identify inaccurate reconstruction without accessing ground-truth information. Based on the study, we propose CoR-GS, which identifies and suppresses inaccurate reconstruction based on the two disagreements: (i) Co-pruning considers Gaussians that exhibit high point disagreement in inaccurate positions and prunes them. (ii) Pseudo-view co-regularization considers pixels that exhibit high rendering disagreement are inaccurate and suppress the disagreement. Results on LLFF, Mip-NeRF360, DTU, and Blender demonstrate that CoR-GS effectively regularizes the scene geometry, reconstructs the compact representations, and achieves state-of-the-art novel view synthesis quality under sparse training views. Project page: https://jiaw-z.github.io/CoR-GS
@@
3d gaussian splatting sparse-view novel view synthesis
@@ 1 Introduction
We hope the observations and discussions in this paper can stimulate further thinking on the randomness of 3D Gaussian radiance fields.
@@ 2 Related Work
2.1 Radiance Fields
Radiance Fields are employed for reconstructing 3D scenes and synthesizing novel views. Neural Radiance Fields (NeRFs) [24] have experienced significant advancements, which learn the neural volumetric representations to represent 3D scenes and render images through volume rendering. Since then, a great number of studies aim to improve the rendering quality [2, 3] and efficiency [5, 25, 32, 11, 43, 21, 13] for NeRFs. Recent advancements in unstructured radiance fields, as demonstrated by studies like [7, 41, 16], employ a collection of Gaussians to represent 3D scenes. Notably, 3D Gaussian Splatting (3DGS) [16] represents scenes using a set of anisotropic 3D Gaussians, and renders images via differentiable splitting. 3DGS has demonstrated remarkable success in reconstructing high-quality complex scenes in real time. Despite the tremendous success of 3DGS in many 3D tasks [22, 38, 33], the behavior of 3DGS with sparse views is less investigated, posing an open problem in the field.
2.2 Novel View Synthesis with Sparse Views
Novel view synthesis aims to generate unseen views of objects or scenes from a set of given images [49, 1]. While given sparse views, many methods tend to reconstruct degraded scenes and render unrealistic novel views. Many studies have explored regularizing NeRFs with sparse views [42, 26, 17, 10]. Some methods focus on designing generative models and pre-train them on extensive datasets [6, 44, 9, 50, 18]. Conversely, others [39, 14, 10, 27, 31, 35] utilize the external knowledge from pre-trained models to regularize the training process.
With the advancements of 3DGS, recent studies [51, 40, 19] have observed the issue of 3DGS to synthesize unrealistic novel views in sparse-view scenarios. These methods rely on predictions of pre-trained depth estimators as regularization to correct the reconstructed geometry. However, external supervision can introduce additional noise, negatively impacting the reconstruction. Our paper introduces a new co-regularization perspective for sparse-view 3DGS by suppressing the disagreement between two 3D Gaussian radiance fields.
2.3 Prediction Agreement
The agreement of predictions of two nerual networks has been explored in many tasks [30, 4, 20, 12, 37, 45, 46], especially for semi-supervised learning [30] and learning with noisy labels [12]. The typical pipeline involves simultaneously training two neural networks. They leverage the agreement of different networks’ predictions to pseudo-label unlabeled data [20] or clean noise from labeled annotations [12], thereby training networks with improved generalization capabilities. Our distinction from previous studies lies in our investigation of disagreement of unstructured 3D Gaussian radiance fields instead of generalized neural networks, where we consider Gaussian positions and rendered results.
@@ 3.1 Preliminary and Definition
Point disagreement. When considering the 3-axis position of Gaussians, Gaussians in two radiance fields can be regarded as two point clouds. We evaluate their differences using Fitness and root mean square error (RMSE), which are used to evaluate the registration between point clouds [48]. Fitness calculates the overlapping area between Gaussians concerning a max distance $\tau=5$ for correspondence. RMSE computes the average distance of correspondent points.
@@ 3.2 Empirical Study
Two 3D Gaussian radiance fields trained with the same sparse views can exhibit different behaviors. We simultaneously train two 3D Gaussian radiance fields and record the disagreements during training. Since ground-truth supervision is directly imposed on training views, we evaluate the rendering disagreement at unseen views. As shown in Fig. 3, the two 3D Gaussian radiance fields exhibit different behaviors. In particular, the two disagreements grow significantly during densification, which creates new Gaussians but locates them blindly to the scene geometry. Therefore, the randomness during densification can be a source of error geometry in sparse-view 3DGS.
@@ 4.1 Co-pruning
The sampling implementation in densification creates new Gaussians blind to the geometry and can be hard to correct with sparse views. Based on the point disagreement, co-pruning identifies and prunes Gaussians that are located in inaccurate positions. We first find the matching correspondence $f:\Theta^{1}\to\Theta^{2}$ based on their nearest Gaussians in the other set:
@@
where $\theta^{1}_{xi},\theta^{1}_{yi},\theta^{1}_{zi}$ represent the 3-axis positions of Gaussians. After the calculation of the non-matching mask for both sets of Gaussians, we prune all Gaussians that are marked as non-matching, which are considered in inaccurate positions. We perform co-pruning every certain number (we set it to 5) of the optimization/density control interleaves, as we aim to prune Gassuains that optimization cannot handle.
@@ 4.2 Pseudo-view Co-regularization
The final training loss is a combination of the ground truth supervised loss at the training view and the color co-regularization term at the pseudo view:
@@ 5.1 Setup
Datasets. We conducted experiments on four datasets: LLFF [23], Mip-NeRF360 [3], DTU [15], and Blender [24]. Our experimental setup aligns with previous studies [26, 42, 35, 51], adopting the same data split, with downsampling rates of $8$, $4$, $4$, and $2$ for LLFF, Mip-NeRF360, DTU, and Blender, respectively. More details are in the Materia. To focus on the target object and eliminate background noise during evaluation, we apply object masks similar to prior works [26] for DTU. Camera poses are assumed to be known based on calibration or other established methods, following the conventions of sparse-view settings.
@@
Implementations. We conduct training for 10k iterations on the LLFF, DTU, and Blender datasets and follow 3DGS [16] to train models for 30k iterations on the MipNeRF dataset. Following FSGS [51], we initialize 3DGS and CoR-GS with fused stereo point cloud from sparse views.
@@ 5.2 Comparison
LLFF. We provide the quantitative results on the LLFF dataset in Tab. 1. Our method consistently achieves the best performance across PSNR, SSIM, LPIPS, and AVGE metrics with 3, 6, and 9 training views. The depth-supervised FSGS [51] enhances the performance of 3DGS in scenarios with 3 and 6 views. However, with 9 training views, 3DGS can already reconstruct the scene structure well, which limits the effectiveness of external depth supervision. Our method conducts co-regularization to avoid reconstructing incorrect geometry, leading to improvements across all metrics with various training views. We provide quantitative visualizations of rendered images and depth at novel views and Gaussian points in Fig. 7. From the rendered depth maps, we observe that vanilla 3DGS can recover certain scene structures but exhibits geometric errors, particularly evident in the unrealistic parts of rendered images. For the point representations, many Gaussians are dispersed throughout space and far from the reconstructed scene. These Gaussians, when observed from novel views, can result in unrealistic rendering images. With depth supervision, FSGS can correct the incorrect geometry of 3DGS; however, noises in external depth maps also adversely affect the geometry. Similarly, relying on depth supervision makes it challenging to directly constrain Gaussians. It can be seen that the Gaussians’ dispersion of FSGS is still not compact. Our method effectively assists Gaussians in reconstructing coherent and compact geometry, resulting in higher-quality novel view rendering.
@@ 5.3 Ablation Study
Co-pruning. Co-pruning improves the geometric structure, reflected in more reasonable depth rendering and more compact Gaussians. However, its drawback is that the rendered depth lacks smoothness, especially evident in the floor part of the horns scene. This is because co-pruning mainly operates on non-matching Gaussians that are far from reconstructed scenes while lacking sufficient constraints on matching Gaussians that are located in reasonable positions.